A chemist wants to mix 10% alcohol solution and a 30% alcohol solution to make a 22% alcohol solution. How many ounces of each solution must be mixed to make 10 ounces of 22% solution

Answer by fractalier(6550)

Let the amount of 10% solution be x. Then 10-x would be the amount of 30% solution. The set up looks like this:
10x + 30(10-x) = 22(10)
10x + 300 - 30x = 220
-20x + 300 = 220
-20x = -80
x = 4 ounces of 10% and then
10-x = 6 ounces of 30%